OpenStudy (anonymous):

Medal and Fan if you can solve this. Point A is translated left 3 units and up 2 units. What is the rule for this translation?

OpenStudy (anonymous):

x' = x-3 y' = y+2 where x,y is start and x',y' is finish

OpenStudy (anonymous):

You win